OpenStudy (anonymous):

9a^-8a-1=(9a+1)( ) how do i figure out whats missing in the parenthesis

OpenStudy (lukebluefive):

For each term on the left side of the expression, ask yourself what should the respective value on the right be multiplied by in order to equal that value? Ignore the middle term for now. For instance, in order for there to be a negative one on the left side of the expression, the positive one on the right side must be multiplied by a negative one. Therefore: \[9a^2 - 8a - 1 = (9a + 1)(? - 1)\] Try to figure out what value goes where "?" is on your own.

OpenStudy (anonymous):

9?

OpenStudy (lukebluefive):

Close. Think of it as the following division problem: \[\frac{ 9a^2 }{ 9a } = ?\]

OpenStudy (anonymous):

is the answer A?

OpenStudy (lukebluefive):

Well, ? = a, so the answer is (a - 1).

OpenStudy (anonymous):

thanks

OpenStudy (lukebluefive):

yw